OpenStudy (snowcrystal):

Please help me! I will fan and medal! Find the discrimination of each quadratic equation.

OpenStudy (snowcrystal):

\[25b^2 - 20b = 0\]

OpenStudy (btaylor):

I think you mean the discriminant. The discriminant is of a polynomial \(ax^2+bx+c\) is found by \(\sqrt{b^2-4ac}\).

OpenStudy (snowcrystal):

It would look like this right??? \[\sqrt{-20^2 - 4* 25}\]

OpenStudy (snowcrystal):

its okay and it would be \[-20^2 -4 * 25 \] then right

OpenStudy (snowcrystal):

im so confused

OpenStudy (snowcrystal):

\[-25 \pm \frac{ \sqrt{65} }{ 2 } \approx -25 \pm \frac{ 8 }{ 2 } = -25\pm 4\]

OpenStudy (anonymous):

Yes. That is correct.

OpenStudy (anonymous):

I am kind of getting confused. I am doing this subject too so I am trying ot help the best way I can.

OpenStudy (snowcrystal):

\[-25 \pm 4 =\] -21 or -29

OpenStudy (anonymous):

Yeah. I think there is supposed to be two numbers at the end.

OpenStudy (anonymous):

Let me double check.

OpenStudy (snowcrystal):

if it is a \[\pm\] it should be

random231 (random231):

 THIS IS NOT THE CORRECT FORMULA!!!